OpenStudy (some.random.cool.kid):

When rounded to the nearest tenth, what are the points of intersection of the system y equals x squared plus three, y equals x plus four?

OpenStudy (some.random.cool.kid):

@aaronq

OpenStudy (some.random.cool.kid):

@StudyGurl14 @ShadowLegendX @sweetburger @johnweldon1993 @jhonyy9 @just_one_last_goodbye @CandyCove

OpenStudy (some.random.cool.kid):

@johnweldon1993

OpenStudy (studygurl14):

You have the system of equations \(\large y=x^2+3\) and \(\large y=x+4\) Set them equal to each other, and solve the quadratic for x

OpenStudy (some.random.cool.kid):

yeah but what about this graph?

OpenStudy (studygurl14):

It's not important. You can solve without the graph

OpenStudy (some.random.cool.kid):

uh ok

OpenStudy (studygurl14):

So, setting them equal to each other, you'd get \(\Large x^2+3=x+4\). Can you simplify that?

OpenStudy (some.random.cool.kid):

well idk i think you can but havent figured out how @StudyGurl14

OpenStudy (some.random.cool.kid):

@jhonyy9 @skullpatrol i need help

OpenStudy (some.random.cool.kid):

@aaronq @agent0smith @Astrophysics